# L’approximation arithmétique de Fujita sur courbes adéliques

### v2

@@ 5.2 Case of $\mathbb{K}$-line bundles
    $\overline{M}$ and $\overline{G}$ are adelic line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, and $X$ be a projective variety over $\operatorname{Spec}K$. Let $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$. Then for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ is effective and
@@
We equip metrics on $L_{1},\ldots,L_{r}$ over $S$, such that $\overline{L}_{1},\ldots,\overline{L}_{r}$ are arithmetically effective. Then for any $b_{1},\ldots,b_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{L}_{1}+\cdots+b_{r}\overline{L}_{r}$ is effective.
By the continuity of $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ from [9, Corollary 6.4.24], for all $\epsilon>0$, there exist $\delta_{1},\ldots,\delta_{r}>0$, such that for all $t_{i}\in(0,\delta_{i})$, $i=1,\ldots,r$, we always have
@@
By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r}\in\mathbb{R}^{+}$ such that $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle. By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is effective. Let
@@
    $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle, and $\overline{G}$ is an arithmetically ample line bundle;
@@
Proof.
By Lemma 5.2.2, for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ is effective and

### v3

@@ 5.2 Case of $\mathbb{K}$-line bundles
    $\overline{M}$ and $\overline{G}$ are adelic line bundles, where $\overline{M}$ has a non-zero small section and $\overline{G}$ is arithmetically ample ;
@@
Let $\overline{L}_{1},\ldots,\overline{L}_{n}$ be a family of adelic line bundle, and $a_{1},\ldots,a_{n}\in\mathbb{R}$. We denote by $\sum\limits_{i=1}^{n}a_{i}L_{i}$ and $\sum\limits_{i=1}^{n}a_{i}\overline{L}_{i}$ the tensor products of line bundles $\bigotimes\limits_{i=1}^{n}L_{i}^{\otimes a_{i}}$ and $\bigotimes\limits_{i=1}^{n}\overline{L}_{i}^{\otimes a_{i}}$ to simplify the notations in the proof below, which are a usual $\mathbb{R}$-line bundle and an adelic $\mathbb{R}$-line bundle respectively.
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, and $X$ be a projective variety over $\operatorname{Spec}K$. Let $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$. Then for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ has a non-zero small section and
@@
We equip metrics on $L_{1},\ldots,L_{r}$ over $S$, such that $\overline{L}_{1},\ldots,\overline{L}_{r}$ have non-zero small sections. Then for any $b_{1},\ldots,b_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{L}_{1}+\cdots+b_{r}\overline{L}_{r}$ has a non-zero small section.
[…]
@@
By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r}\in\mathbb{R}^{+}$ such that $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle. By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ has a non-zero small section. Let
@@
    $\overline{M}$ is an $\mathbb{R}$-line bundle having a non-zero small section, and $\overline{G}$ is an arithmetically ample line bundle ;
@@
Démonstration.
By Lemma 5.2.2, for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ has a non-zero small section and
@@
Declaration
Conflict of interest statement
The author declares that there is no potential conflict of interest.
Ethical statement
The author guarantees that the manuscript has neither been published elsewhere nor submitted for publication in another journal, and will not be submitted to other journals while being reviewed. The author certify that they have written entirely the manuscript and that its main results are original, and cite all necessary sources. The author will inform immediately to the editors if any error is found in the manuscript after the submission.
The data availability statement
The data that support the findings of this study are openly available.
Références
  • [1] J.-B. Bost & H. Chen – « Concerning the semistability of tensor products in Arakelov geometry », Journal de Mathématiques Pures et Appliquées. Neuvième Série 99 (2013), no. 4, p. 436–488.
  • [2] S. Boucksom & H. Chen – « Okounkov bodies of filtered linear series », Compositio Mathematica 147 (2011), no. 4, p. 1205–1229.
  • [3] N. BourbakiÉléments de mathématique. Fasc. XIII. Livre VI : Intégration. Chapitres 1, 2, 3 et 4 : Inégalités de convexité, Espaces de Riesz, Mesures sur les espaces localement compacts, Prolongement d’une mesure, Espaces $L\mathrm{sp}{p}$, Deuxième édition revue et augmentée. Actualités Scientifiques et Industrielles, No. 1175, Hermann, Paris, 1965.
  • [4] H. Chen – « Arithmetic Fujita approximation », Annales Scientifiques de l’École Normale Supérieure. Quatrième Série 43 (2010), no. 4, p. 555–578.
  • [5] — , « Convergence des polygones de Harder-Narasimhan », Mémoires de la Société Mathématique de France 120 (2010), p. 1–120.
  • [6] — , « Differentiability of the arithmetic volume function », Journal of the London Mathematical Society. Second Series 84 (2011), no. 2, p. 365–384.
  • [7] — , « Newton-Okounkov bodies : an approach of function field arithmetic », Journal de Théorie des Nombres de Bordeaux 30 (2018), no. 3, p. 829–845.
  • [8] H. Chen & H. Ikoma – « On subfiniteness of graded linear series », European Journal of Mathematics 6 (2020), no. 2, p. 367–399.
  • [9] H. Chen & A. MoriwakiArakelov geometry over adelic curves, Lecture Notes in Mathematics, vol. 2258, Springer, Singapore, [2020] ©2020.
  • [10] — , « Arithmetic intersection theory over adelic curves », (2021), to appear in Mémoires de la Société Mathématique de France. Nouvelle Série.
  • [11] — , Positivity in Arakelov geometry over adelic curves—Hilbert-Samuel formula and equidistribution theorem, Progress in Mathematics, vol. 355, Birkhäuser/Springer, Cham, [2024] ©2024.
  • [12] T. Fujita – « Approximating Zariski decomposition of big line bundles », Kodai Mathematical Journal 17 (1994), no. 1, p. 1–3.
  • [13] G. Harder & M. S. Narasimhan – « On the cohomology groups of moduli spaces of vector bundles on curves », Mathematische Annalen 212 (1974/1975), p. 215–248.
  • [14] H. Ikoma – « On the concavity of the arithmetic volumes », International Mathematics Research Notices. IMRN (2015), no. 16, p. 7063–7109.
  • [15] — , « Adelic Cartier divisors with base conditions and the continuity of volumes », Kyoto Journal of Mathematics 61 (2021), no. 4, p. 905–947.
  • [16] R. LazarsfeldPositivity in algebraic geometry. I, Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. A Series of Modern Surveys in Mathematics, vol. 48, Springer-Verlag, Berlin, 2004, Classical setting : line bundles and linear series.
  • [17] — , Positivity in algebraic geometry. II, Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. A Series of Modern Surveys in Mathematics, vol. 49, Springer-Verlag, Berlin, 2004, Positivity for vector bundles, and multiplier ideals.
  • [18] R. Lazarsfeld & M. Mustaţă – « Convex bodies associated to linear series », Annales Scientifiques de l’École Normale Supérieure. Quatrième Série 42 (2009), no. 5, p. 783–835.
  • [19] A. Moriwaki – « Arithmetic height functions over finitely generated fields », Inventiones Mathematicae 140 (2000), no. 1, p. 101–142.
  • [20] — , « Continuity of volumes on arithmetic varieties », Journal of algebraic geometry 18 (2009), no. 3, p. 407–457.
  • [21] — , « Continuous extension of arithmetic volumes », International Mathematics Research Notices. IMRN (2009), no. 19, p. 3598–3638.
  • [22] — , « Adelic divisors on arithmetic varieties », Memoirs of the American Mathematical Society 242 (2016), no. 1144, p. v+122.
  • [23] A. Sédillot – « Differentiability of the $\chi$-volume function over an adelic curve », (2023), arXiv:2303.03377.
  • [24] S. Takagi – « Fujita’s approximation theorem in positive characteristics », Journal of Mathematics of Kyoto University 47 (2007), no. 1, p. 179–202.
  • [25] X. Yuan – « Big line bundles over arithmetic varieties », Inventiones Mathematicae 173 (2008), no. 3, p. 603–649.
  • [26] — , « On volumes of arithmetic line bundles », Compositio Mathematica 145 (2009), no. 6, p. 1447–1464.
  • [27] — , « Volumes of arithmetic Okounkov bodies », Mathematische Zeitschrift 280 (2015), no. 3-4, p. 1075–1084.
  • [28] O. Zariski – « The theorem of Riemann-Roch for high multiples of an effective divisor on an algebraic surface », Annals of Mathematics, Second Series 76 (1962), p. 560–615.
  • [29] S. Zhang – « Positive line bundles on arithmetic varieties », Journal of the American Mathematical Society 8 (1995), no. 1, p. 187–221.